OpenStudy (anonymous):

Give an expression. INVENTORY -- The school kitchen had 11 cases of juice plus 3 extra cans of juice. After lunch they had 6 cases of juice and no extra cans. How many cans of juice were sold during lunch? (Note: A case contains 24 cans.)

OpenStudy (agreene):

24*11+3 = number of cans before lunch (267) 24*6 = number of cans after lunch (144) number before number after = number sold 267-144 = 123 cans of juice were sold during lunch.